OpenStudy (anonymous):

The length of the diameter of a sphere with a volume of 321,57 miles cubed is 4.25 miles. TRUE or FALSE ?

OpenStudy (anonymous):

Use the formula : \[Volume = 4\pi r^2/3\]

OpenStudy (anonymous):

I DONT KNOW THE DIAMETER SO I DONT KNOW WHAT THE RADIUS IS TO SOLVE THIS.

OpenStudy (anonymous):

never mmind lol. hold on

OpenStudy (anonymous):

\[Radius = \sqrt{3*Volume/4\pi}\]

OpenStudy (anonymous):

so is the answer FALSE ?

OpenStudy (anonymous):

Yes it is...